OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). tan x + sec x = 1

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Change everything to sine and cosine

OpenStudy (anonymous):

im not sure how to do that @Mertsj

OpenStudy (mertsj):

What is the definition of tangent?

OpenStudy (anonymous):

opposite/adjacent

OpenStudy (mertsj):

In terms of sin and cosine?

OpenStudy (anonymous):

sin x /cos x

OpenStudy (mertsj):

yes. And what is sec?

OpenStudy (anonymous):

1/cos x

OpenStudy (mertsj):

Yes. So now we have:

OpenStudy (anonymous):

sin x /cos x + 1/ cos x =1 ?

OpenStudy (mertsj):

\[\frac{\sin x}{cosx}+\frac{1}{\cos x}=1\]

OpenStudy (mertsj):

So multiply each term by cos x

OpenStudy (anonymous):

sinxcosx + cosx = cosx

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

\[\frac{\sin x}{\cos x}\times cosx +\frac{1}\cos x \times \cos x=1\times \cos x\]

OpenStudy (anonymous):

sin x + cos x = cos x ?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

sinx + 1 = cosx

OpenStudy (anonymous):

now what do i do?

OpenStudy (anonymous):

x = x = 0 x = No solution

OpenStudy (anonymous):

x = π/4 x = 0 x = 5π/4 No solution

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

no solution

OpenStudy (anonymous):

thank u

OpenStudy (mertsj):

No wait. I think that 0 might work.

OpenStudy (mertsj):

Try it

OpenStudy (anonymous):

i did. its 0. @Mertsj

OpenStudy (mertsj):

Good.